# Risk-Averse Total-Reward Reinforcement Learning

### v1

@@ 12 $x_{\max}\leftarrow\max\tilde{x}(s,a,\beta_{c}),s\in\mathcal{S},a\in\mathcal{A}$ ;
From line 4 of Algorithm 1, we have
@@
where $\|r\|_{\infty}=\max_{s,s^{\prime}\in\mathcal{S},a\in\mathcal{A}}|r(s,a,s^{\prime})|$ and $|q|_{\max}=\max_{s\in\mathcal{S},a\in\mathcal{A}}|q_{i}(s,a,\beta)|,i\in\mathbb{N}$.
Let us estimate $|q|_{\max}$. For any random variable $\tilde{x}$, for any risk level $\beta\in\mathcal{B}$, $\operatorname{ERM}_{\beta}\left[\tilde{x}\right]$ is non-increasing and satisfies monotonicity [19, Lemma A.7 and Lemma A.8],
@@
$=\max\{|\mathbb{E}[\tilde{x}]-\beta(x_{\max}-x_{\min})^{2}/8|,|\mathbb{E}[\tilde{x}]|\}$
As mentioned in Section 2,
@@
We estimate $\mathbb{E}[\tilde{x}]$, $x_{\max}$, $x_{\min}$ by setting $\beta_{c}=1^{-10}$ in Algorithm 3. Therefore, for any $\beta\in\mathcal{B}$,
$z_{\min}(\beta)=-||r||_{\infty}-2\max\{|c-\beta\cdot d|,|c|\},\quad z_{\max}(\beta)=||r||_{\infty}+2\max\{|c-\beta\cdot d|,|c|\}$
@@ A.4 $\beta_{0}$ Derivation
For a random variable $\tilde{x}$ and $\beta>0$ [19, Lemma A.8]
@@
where $x_{\max}$ and $x_{\min}$ are estimated in Algorithm 3.
@@ B.1 Standard Convergence Results
[…]
@@
Note that B.2 leads to a convergence result somewhat weaker than the results for weighted maximum norm pseudo-contraction. Then we need a separate boundedness condition [8, proposition 4.6], which is restated here as Proposition B.3.
@@
Let $r_{t}$ be the sequence generated by the iteration shown in Equation 21. We assume that the $b$th component $\tilde{r}(b)$ of $\tilde{r}$ is updated according to Equation 21, with the understanding that $\tilde{\theta}_{i}(b)=0$ if $\tilde{r}(b)$ is not updated at iteration $i$. $\tilde{\phi}_{i}(b)$ is a random noise term. Then we assume the following:
@@
    The noise terms $\tilde{\phi}_{i}(b)$ satisfies B.1,
@@ B.2.1 Operator Definitions
Let us define some operators $G$ and $H$ for the convergence analysis of Algorithm 1 and Algorithm 2. Let $b=(s,a,\beta),s\in\mathcal{S},a\in\mathcal{A}$, $\beta\in\mathcal{B}$, and $\xi>0$. $(Gq)(b)$ can be interpreted as the average gradient steps. $(G_{s^{\prime}}q)(b)$ can be interpreted as the individual gradient step for a specific next state $s^{\prime}$.
$(Gq)(b)$ $:=\frac{\partial}{\partial y}\mathbb{E}^{a,s}\Bigl{[}\ell_{\beta}\Bigl{(}r(s,a,\tilde{s}_{1})+\max_{a^{\prime}\in\mathcal{A}}q(\tilde{s}_{1},a^{\prime},\beta)-y\Bigr{)}\Bigr{]}\mid_{y=q(s,a,\beta)}$ (22)
@@
$(G_{s^{\prime}}q)(b)$ $:=\frac{\partial}{\partial y}\mathbb{E}\Bigl{[}\ell_{\beta}\Bigl{(}r(s,a,s^{\prime})+\max_{a^{\prime}\in\mathcal{A}}q(s^{\prime},a^{\prime},\beta)-y\Bigr{)}\Bigr{]}\mid_{y=q(s,a,\beta)}$ (23)
@@
$(Hq)(b)$ can be interpreted as the $q$ update for a random variable representing the next state. $(H_{s^{\prime}}q)(b)$ can be interpreted as the $q$ update a specific next state $s^{\prime}$.
$(Hq)(b):=q(b)-\xi\cdot(Gq)(b),\qquad(H_{s^{\prime}}q)(b):=q(b)-\xi\cdot(G_{s^{\prime}}q)(b)$ (24)
Consider a random sequence of inputs $(\tilde{s}_{i},\tilde{a}_{i})_{i=0}^{\infty}$ in Algorithm 1. We can define a real-valued random variable $\tilde{\phi}(s,a)$ for $s\in\mathcal{S},a\in\mathcal{A},\beta>0$ in (25).
$\tilde{\phi}_{i}(s,a,\beta)=\begin{cases}(H_{\tilde{s}_{i}}\tilde{q}_{i})(s,a)-(H\tilde{q}_{i})(s,a),&\text{if }(\tilde{s}_{i},\tilde{a}_{i})=(s,a)\\ 0,&\text{otherwise }\\ \end{cases}$ (25)
@@
$\tilde{\theta}_{i}(s,a,\beta)=\begin{cases}\tilde{\eta}_{i}/\xi,&\text{if }(\tilde{s}_{i},\tilde{a}_{i})=(s,a)\\ 0,&\text{otherwise }\\ \end{cases}$ (26)
@@
The random sequences of iterations followed by Algorithm 1 satisfies
@@
where the terms are defined in Equations 22, 23, 24, 25 and 26.
@@
Case 2: suppose that $\tilde{b}=(\tilde{s},\tilde{a},\beta)\neq(s,a,\beta)=b$, then by algebraic manipulation, the algorithm does not change the $q$-function:
@@ B.2.2 Random Noise Analysis
We restate Lemma C.16 in [17] here as Lemma B.5, which is useful for proving properties of the random noise.
@@
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\xi\cdot\mathbb{E}[\mathbb{E}^{a,s}[-(G_{\tilde{s}_{1}}\tilde{q}_{i})(b)]+(G\tilde{q}_{i})(b)\mid\mathcal{F}_{i},\tilde{b}_{i}=b]$
@@
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}$ $\xi^{2}\cdot\mathbb{E}\Biggl{[}\mathbb{E}^{a,s}\Bigl{[}\Bigl{(}(G_{\tilde{s}_{1}}\tilde{q}_{i})(b)-(G\tilde{q}_{i})(b)\Bigr{)}^{2}\Bigr{]}\mid\mathcal{F}_{i},\tilde{b}_{i}=b\Biggr{]}$
=$\stackrel{{\scriptstyle\text{(b)}}}{{=}}$ $\xi^{2}\cdot\mathbb{E}\biggl{[}\mathbb{E}^{a,s}\Bigl{[}\Bigl{(}\mathbb{E}[\partial\ell_{\beta}(\tilde{\delta_{i}}(\tilde{s}_{1},\beta))\mid\tilde{s}_{1}]-\mathbb{E}^{a,s}[\partial\ell_{\beta}(\tilde{\delta_{i}}(\tilde{s}_{1},\beta))]\Bigr{)}^{2}\Bigr{]}\mid\mathcal{F}_{i},\tilde{b}_{i}=b\biggr{]}$
=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}$ $\xi^{2}\cdot\mathbb{E}\biggl{[}\Bigl{(}\mathbb{E}^{a,s}\Bigl{[}\bigl{(}\mathbb{E}[\partial\ell_{\beta}(\tilde{\delta_{i}}(\tilde{s}_{1},\beta))\mid\tilde{s}_{1}]\bigr{)}^{2}\Bigr{]}-\Bigl{(}\mathbb{E}^{a,s}[\partial\ell_{\beta}(\tilde{\delta_{i}}(\tilde{s}_{1},\beta))]\Bigr{)}^{2}\Bigr{]}\Bigr{)}\mid\mathcal{F}_{i},\tilde{b}_{i}=b\biggr{]}$
@@
Step $(a)$ follows Lemma B.5 given that the randomness of $-(G_{\tilde{s}^{\prime}_{i}}\tilde{q}_{i})(b)+(G\tilde{q}_{i})(b)$ only comes from $\tilde{s}^{\prime}_{i}$ when conditioning on $\mathcal{F}_{i},\tilde{b}_{i}=b,\tilde{\eta}_{i}$ and $\mathcal{G}_{i-1}$. Step $(b)$ follows by substituting $(G_{\tilde{s}^{\prime}_{i}}\tilde{q}_{i})(b)$ with (22), substituting $(G\tilde{q}_{i})(b)$ with (23), and replacing by using $\tilde{\delta}_{i}$ defined in (29). The equality in step $(c)$ holds because for a random variable $\tilde{x}=\mathbb{E}[\partial\ell_{\beta}(\tilde{\delta_{i}}(\tilde{s}_{1},\beta))\mid\tilde{s}_{1}]$, the variance satisfies $\mathbb{E}[(\tilde{x}-\mathbb{E}[\tilde{x}])^{2}]=\mathbb{E}[\tilde{x}^{2}]-(\mathbb{E}[\tilde{x}])^{2}$. Step $(d)$ upper bounds the expectation by a maximum. Step $(e)$ uses $\partial\ell_{\beta}(0)=0$ from the definition in (9). Step $(f)$ uses Lemma 4.4 to bound the derivative difference as a function of the step size. Step $(g)$ derives the final upper bound since
@@ B.2.3 Monotonicity of Operator $H$
Let us restate [17, C.13] as Lemma B.8, which is useful to prove that the operator $H$ satisfies monotonicity.
Lemma B.8 (C.13 in [17]).
Suppose that $f:\mathbb{R}\rightarrow\mathbb{R}$ is a differentiable $\mu-$strongly convex function with an $L-$Lipschitz continuous gradient. Consider $x_{i}\in\mathbb{R}$ and a gradient update for any step size $\xi\in(0,1/L]:$
$x_{i+1}:=x_{i}-\xi\cdot f^{\prime}(x_{i})$
Then $\exists l\in[1/L,1/\mu]$ such that $\xi/l\in(0,1]$ and
$x_{i+1}=(1-\xi/l)\cdot x_{i}+\xi/l\cdot x^{\star}$
where $x^{\star}=\operatorname*{arg\,min}_{x\in\mathbb{R}}f(x)$ (unique from strong convexity)
Lemma B.9.
$x(b)\leq y(b)$ $\Rightarrow(Hx)(b)\leq(Hy)(b)$ (a)
$(Hq^{\star})(b)$ $=q^{\star}(b)$ (b)
Proof.
@@
The function $f$ is strongly convex with Lipschitz gradient with parameters $\ell$ and $L$ based on Lemma 4.4. Let $y^{\star}=\arg\min_{y\in\mathbb{R}}f(y)$ and $\exists l\in[1/L,1/\ell$] such that
@@
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}q(b)-\xi f^{\prime}(q(b))$
=$\stackrel{{\scriptstyle\text{(b)}}}{{=}}(1-\xi/l)q(b)+\xi/l\cdot y^{\star}$
=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}(1-\xi/l)q(b)+\xi/l\cdot(\hat{B}q)(b)$
Step (a) follows the replacement with (22) and (31). Step(b) follows the Lemma B.8. Step(c) follows the fact that $\hat{B}$ is the ERM Bellman operator defined in (11) and $y^{\star}$ is the unique solution to (10).
@@
From [43, theorem 3.3], we know that $q^{\star}(b)$ is a fixed point of $\hat{B}$. Then $q^{\star}(b)$ is also a fixed point of $H$. That is, $(Hq^{\star})(b)=q^{\star}(b)$.
@@
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}$ $(1-\xi/l)(q-c\cdot e)+(\xi/l)(\hat{B}(q-c\cdot e))$
@@
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}$ $(1-\xi/l)(q+c\cdot e)+(\xi/l)\hat{B}(q)+(\xi/l)\cdot c\cdot e$
@@
Proof of Theorem 4.2 We verify that the sequence of our $q$-learning iterates satisfies the properties in Proposition B.3. The step size condition in Theorem 4.2 guarantees that we satisfy property $(a)$ in Proposition B.3. Lemma B.6 and Lemma B.7 show that we satisfy property (b) in Proposition B.3. Lemma B.9 shows that we satisfy property $(c)$ in Proposition B.3. ∎
@@
First, from Theorem 3.3, we have
@@
For the part $(b)$, $\ell_{\beta}$ is a strongly convex function because $\ell^{\prime\prime}_{\beta}(y)=\beta\cdot e^{-\beta\cdot(\tilde{x}-y)}>0$. Then there will be a unique $y^{\star}$ value such that $\mathbb{E}^{a,s}\Bigl{[}\ell_{\beta}\bigl{(}r(s,a,\tilde{s}_{1})+\max_{a^{\prime}\in\mathcal{A}}q(\tilde{s}_{1},a^{\prime},\beta)-y\bigr{)}\Bigr{]}$ attains the minimum value, and the $y^{\star}$ value is equal to $q^{\star}(s,a,\beta)$, which follows from Proposition 3.2. So $\hat{B}_{\beta}q^{\star}=q^{\star}$
@@
From Theorem 4.2, for some $\beta\in\mathcal{B}$, ERM-Q learning algorithm converges to the optimal ERM value function. Theorem 3.4 shows that there exists $\delta-$optimal policy such that it is an ERM-TRC optimal for some $\beta\in\mathcal{B}$. It is sufficient to compute an ERM-TRC optimal policy for one of those $\beta$ values. This analysis shows that Algorithm 2 converges to its optimal EVaR value funciton. ∎
Appendix C Additional Material of Section 5
The machine used to conduct all experiments referenced in Section 5 is a single machine with the following specifications:

### v2

@@ 12 $x_{\max}\leftarrow\max\tilde{x}(s,a,\beta_{c}),s\in\mathcal{S},a\in\mathcal{A}$ ;
From line 4 of Algorithm˜1, we have
@@
where $\|r\|_{\infty}=\max_{s,s^{\prime}\in\mathcal{S},a\in\mathcal{A}}|r(s,a,s^{\prime})|$ and $|q|_{\max}=\max_{s\in\mathcal{S},a\in\mathcal{A}}|\tilde{q}_{i}(s,a,\beta)|,i\in\mathbb{N}$.
Let us estimate $|q|_{\max}$. For any random variable $\tilde{x}$, for any risk level $\beta\in\mathcal{B}$, $\operatorname{ERM}_{\beta}\left[\tilde{x}\right$] is non-increasing in Lemma˜A.2 and satisfies monotonicity in Lemma˜A.3,
@@
[…]
As mentioned in Section˜2,
@@
We estimate $\mathbb{E}[\tilde{x}$], $x_{\max}$, $x_{\min}$ by setting $\beta_{c}=1^{-10}$ in Algorithm˜3. Therefore, for any $\beta\in\mathcal{B}$,
[…]
@@ A.5 $\beta_{0}$ Derivation
From Lemma˜A.3, for a random variable $\tilde{x}$ and $\beta>0$, we have
@@
where $x_{\max}$ and $x_{\min}$ are estimated in Algorithm˜3.
@@ B.1 Standard Convergence Results
    There exists a norm $\|\cdot\|$ on $\mathbb{R}^{\mathcal{N}}$ and constants $A$ and $B$ such that
@@
Note that ˜B.2 leads to a convergence result somewhat weaker than the results for weighted maximum norm pseudo-contraction. Then we need a separate boundedness condition [8, proposition 4.6], which is restated here as Proposition˜B.3.
@@
Let $r_{t}$ be the sequence generated by the iteration shown in Equation˜26. We assume that the $b$th component $\tilde{r}(b)$ of $\tilde{r}$ is updated according to Equation˜26, with the understanding that $\tilde{\theta}_{i}(b)=0$ if $\tilde{r}(b)$ is not updated at iteration $i$. $\tilde{\phi}_{i}(b)$ is a random noise term. Then we assume the following:
@@
    The noise terms $\tilde{\phi}_{i}(b)$ satisfies ˜B.1,
@@ B.2.1 Operator Definitions
Let $b=(s,a,\beta),s\in\mathcal{S},a\in\mathcal{A}$, $\beta\in\mathcal{B}$, and $\xi>0$, we use some operators $G$ defined in (12) and (13), and $H$ defined in (14) for the convergence analysis of Algorithm˜1 and Algorithm˜2.
$(Gq)(b)$ $=\frac{\partial}{\partial y}\mathbb{E}^{a,s}\Bigl[\ell_{\beta}\Bigl(r(s,a,\tilde{s}_{1})+\max_{a^{\prime}\in\mathcal{A}}q(\tilde{s}_{1},a^{\prime},\beta)-y\Bigr)\Bigr]\mid_{y=q(s,a,\beta)}$
@@
$(G_{s^{\prime}}q)(b)$ $=\frac{\partial}{\partial y}\mathbb{E}\Bigl[\ell_{\beta}\Bigl(r(s,a,s^{\prime})+\max_{a^{\prime}\in\mathcal{A}}q(s^{\prime},a^{\prime},\beta)-y\Bigr)\Bigr]\mid_{y=q(s,a,\beta)}$
@@
$(Hq)(b):=q(b)-\xi\cdot(Gq)(b),\qquad(H_{s^{\prime}}q)(b):=q(b)-\xi\cdot(G_{s^{\prime}}q)(b)$
Consider a random sequence of inputs $(\tilde{s}_{i},\tilde{a}_{i})_{i=0}^{\infty}$ in Algorithm˜1. We can define a real-valued random variable $\tilde{\phi}(s,a)$ for $s\in\mathcal{S},a\in\mathcal{A},\beta>0$ in (27).
[…]
@@
[…]
@@
The random sequences of iterations followed by Algorithm˜1 satisfies
@@
where the terms are defined in Equations˜12, 13, 14, 27 and 28.
@@
Case 2: suppose that $\tilde{b}=(\tilde{s},\tilde{a},\beta)\neq(s,a,\beta)=b$, then by algebraic manipulation, the algorithm does not change the $q$:
@@ B.2.2 Random Noise Analysis
We restate Lemma C.16 in [17] here as Lemma˜B.5, which is useful for proving properties of the random noise.
@@
$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\xi\cdot\mathbb{E}[\mathbb{E}^{a,s}[-(G_{\tilde{s}_{1}}\tilde{q}_{i})(b)]+(G\tilde{q}_{i})(b)\mid\mathcal{F}_{i},\tilde{b}_{i}=b]$
@@
$\stackrel{{\scriptstyle\text{(a)}}}{{=}}$ $\xi^{2}\cdot\mathbb{E}\Biggl[\mathbb{E}^{a,s}\Bigl[\Bigl((G_{\tilde{s}_{1}}\tilde{q}_{i})(b)-(G\tilde{q}_{i})(b)\Bigr)^{2}\Bigr]\mid\mathcal{F}_{i},\tilde{b}_{i}=b\Biggr]$
$\stackrel{{\scriptstyle\text{(b)}}}{{=}}$ $\xi^{2}\cdot\mathbb{E}\biggl[\mathbb{E}^{a,s}\Bigl[\Bigl(\mathbb{E}[\partial\ell_{\beta}(\tilde{\delta_{i}}(\tilde{s}_{1},\beta))\mid\tilde{s}_{1}]-\mathbb{E}^{a,s}[\partial\ell_{\beta}(\tilde{\delta_{i}}(\tilde{s}_{1},\beta))]\Bigr)^{2}\Bigr]\mid\mathcal{F}_{i},\tilde{b}_{i}=b\biggr]$
$\stackrel{{\scriptstyle\text{(c)}}}{{=}}$ $\xi^{2}\cdot\mathbb{E}\biggl[\Bigl(\mathbb{E}^{a,s}\Bigl[\bigl(\mathbb{E}[\partial\ell_{\beta}(\tilde{\delta_{i}}(\tilde{s}_{1},\beta))\mid\tilde{s}_{1}]\bigr)^{2}\Bigr]-\Bigl(\mathbb{E}^{a,s}[\partial\ell_{\beta}(\tilde{\delta_{i}}(\tilde{s}_{1},\beta))]\Bigr)^{2}\Bigr]\Bigr)\mid\mathcal{F}_{i},\tilde{b}_{i}=b\biggr]$
@@
Step $(a)$ follows Lemma˜B.5 given that the randomness of $(G_{\tilde{s}^{\prime}_{i}}\tilde{q}_{i})(b)+(G\tilde{q}_{i})(b)$ only comes from $\tilde{s}^{\prime}_{i}$ when conditioning on $\mathcal{F}_{i},\tilde{b}_{i}=b,\tilde{\eta}_{i}$ and $\mathcal{G}_{i-1}$. Step $(b)$ follows by substituting $(G_{\tilde{s}^{\prime}_{i}}\tilde{q}_{i})(b)$ with (12), substituting $(G\tilde{q}_{i})(b)$ with (13), and replacing by using $\tilde{\delta}_{i}$ defined in (31). The equality in step $(c)$ holds because for a random variable $\tilde{x}=\mathbb{E}[\partial\ell_{\beta}(\tilde{\delta_{i}}(\tilde{s}_{1},\beta))\mid\tilde{s}_{1}]$, the variance satisfies $\mathbb{E}[(\tilde{x}-\mathbb{E}[\tilde{x}])^{2}]=\mathbb{E}[\tilde{x}^{2}]-(\mathbb{E}[\tilde{x}])^{2}$. Step $(d)$ upper bounds the expectation by a maximum. Step $(e)$ uses $\partial\ell_{\beta}(0)=0$ from the definition in (9). Step $(f)$ uses Lemma˜4.4 to bound the derivative difference as a function of the step size. Step $(g)$ derives the final upper bound since
@@ B.2.3 Monotonicity of Operator $H$
Let us prove that the operator $H$ defined in (14) satisfies monotonicity.
Proof of Lemma˜3.4.
@@
The function $f$ is strongly convex with a Lipschitz-continuous gradient with parameters $\ell$ and $L$ based on Lemma˜4.4. Let $y^{\star}=\arg\min_{y\in\mathbb{R}}f(y)$ and $\exists l\in[1/L,1/\ell$] such that
@@
$\stackrel{{\scriptstyle\text{(a)}}}{{=}}q(b)-\xi f^{\prime}(q(b))$
$\stackrel{{\scriptstyle\text{(b)}}}{{=}}(1-\xi/l)q(b)+\xi/l\cdot y^{\star}$
$\stackrel{{\scriptstyle\text{(c)}}}{{=}}(1-\xi/l)q(b)+\xi/l\cdot(\hat{B}q)(b)$
Step (a) follows the replacement with (12) and (33). Step(b) follows the Lemma˜2.1. Step(c) follows the fact that $\hat{B}$ is the ERM Bellman operator defined in (11) and $y^{\star}$ is the unique solution to (10).
@@
From [44, theorem 3.3], we know that $q^{\star}(b)$ is a fixed point of $\hat{B}$. Then $q^{\star}(b)$ is also a fixed point of $H$. That is, $(Hq^{\star})(b)=q^{\star}(b)$.
@@
$\stackrel{{\scriptstyle\text{(a)}}}{{=}}$ $(1-\xi/l)(q-c\cdot e)+(\xi/l)(\hat{B}(q-c\cdot e))$
@@
$\stackrel{{\scriptstyle\text{(a)}}}{{=}}$ $(1-\xi/l)(q+c\cdot e)+(\xi/l)\hat{B}(q)+(\xi/l)\cdot c\cdot e$
@@
Proof of Theorem˜4.2 We verify that the sequence of our Q-learning iterates satisfies the properties in Proposition˜B.3. The step size condition in Theorem˜4.2 guarantees that we satisfy property $(a)$ in Proposition˜B.3. Lemma˜B.6 and Lemma˜B.7 show that we satisfy property (b) in Proposition˜B.3. Lemma˜3.4 shows that we satisfy property $(c)$ in Proposition˜B.3. ∎
@@
First, from Theorem˜3.3, we have
@@
For the part $(b)$, $\ell_{\beta}$ is a strongly convex function because $\ell^{\prime\prime}_{\beta}(y)=\beta\cdot e^{-\beta\cdot(\tilde{x}-y)}>0$. Then there will be a unique $y^{\star}$ value such that $\mathbb{E}^{a,s}\Bigl[\ell_{\beta}\bigl(r(s,a,\tilde{s}_{1})+\max_{a^{\prime}\in\mathcal{A}}q(\tilde{s}_{1},a^{\prime},\beta)-y\bigr)\Bigr]$ attains the minimum value, and the $y^{\star}$ value is equal to $q^{\star}(s,a,\beta)$, which follows from Proposition˜3.2. So $\hat{B}_{\beta}q^{\star}=q^{\star}$
@@
From Theorem˜4.2, for some $\beta\in\mathcal{B}$, ERM-Q learning algorithm converges to the optimal ERM value function. Theorem˜3.5 shows that there exists $\delta$-optimal policy such that it is an ERM-TRC optimal for some $\beta\in\mathcal{B}$. It is sufficient to compute an ERM-TRC optimal policy for one of those $\beta$ values. This analysis shows that Algorithm˜2 converges to its optimal EVaR value funciton. ∎
Appendix C Additional Material of Section˜5
The machine used to conduct all experiments referenced in Section˜5 is a single machine with the following specifications: